OpenStudy (anonymous):

x=-5+4i (i=-1^1/2) Calculate the value of x^4+9x^3+35x^2-x+4...........reply soon

OpenStudy (anonymous):

have answer chocies

OpenStudy (anonymous):

nopes....it is subjective

OpenStudy (anonymous):

Could you show us what you've done so far please?

OpenStudy (anonymous):

yea wat azteck said

OpenStudy (anonymous):

x+5=4i x^2+10x+41=0 then, x^2(x^2+10x+41)-x^3-6x^2-x+4 I dont know wt to do beyond this........

OpenStudy (anonymous):

Not quite. What's \[\large -1^{\frac{1}{2}}=?\]

OpenStudy (anonymous):

You wrote this i=-1^1/2 . Could you draw it using the drawing tool, because there's quite a lot of expressions that can be made when looking at that.

OpenStudy (anonymous):

\[\sqrt{-1}\] it is related to complex no.s

OpenStudy (anonymous):

it is called iota and is represented as i

OpenStudy (anonymous):

Ah okay. Yep I see. I'm not sure, but I think it could be good to use long division.

OpenStudy (anonymous):

You know x=-5+4i is a root. So (x+5-4i) is a factor of the polynomial.

OpenStudy (anonymous):

x^2+10x+41=0 How did you get that?

OpenStudy (anonymous):

alright let me check if i am getting the answer

OpenStudy (anonymous):

ah I see, you squared both sides.

OpenStudy (anonymous):

don't think that's gonna work for you. It's best to do long diviosnl

OpenStudy (anonymous):

division*

OpenStudy (anonymous):

I may be able to use wat you've done so far. You have this right? \[\large x^2+10x+41=0\] SO we make x^2 the subject. \[\large x^2=-10x-41\] \[x^4+9x^3+35x^2-x+4= (-10x-41)^2+9x(-10x-41)+35(-10x-41)-x+4\]

OpenStudy (anonymous):

\[=100x^2+820x+1681-90x-369x-350x-1435-x+4\] \[=100x^2+10x+250\] \[=100(-10x-41)+10(-5+4i)+250\] \[=-100x-4100-50+40i+250\] \[=-100(-5+4i)-3900+40i\] \[=500-400i-3900+40i\] \[=-3400-360i\] You can use de Moivre's theorem, but it won't give you a a precise angle. Long division wouldn't work as well as I thought it would.

OpenStudy (anonymous):

But what you did was a good start.

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

@Azteck i will tell you something tht long division works better. first squaring as in x^2 + 10x + 41 Then dividing x^4+9x^3+35x^2-x+4 and u get the quotient as x^2-x+4 and remainder -160. we know that x^2+10x+41 = 0. So when we multiply Quotient with divisor we get 0. and the remainder -160 is the answer.

OpenStudy (anonymous):

@Azteck Thanks for ur help anyways

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

@snowman97 I said thanks to @Azteck

OpenStudy (anonymous):

want to message each other

OpenStudy (anonymous):

@snowman97 Abt wt? and why?

OpenStudy (anonymous):

just talk

OpenStudy (anonymous):

Now worries.

OpenStudy (anonymous):

no*

OpenStudy (anonymous):

hmmm